OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9>

sam (.sam.):

\[\cos \theta = \frac{u \cdot v}{|u||v|}\]

sam (.sam.):

For dot product